OpenStudy (anonymous):

Indicate in standard form the equation of the line through the given points. P(0, -4), Q(5, 1)

OpenStudy (anonymous):

@Hero i need help

OpenStudy (anonymous):

URGENT!!!!!

OpenStudy (richyw):

well you can get the slope easily

OpenStudy (anonymous):

but the standard form

OpenStudy (richyw):

and then you already know the y intercept from the first coordinate given

OpenStudy (anonymous):

can you help pleaze

OpenStudy (saifoo.khan):

For urgent help, please contact 911.

hero (hero):

@777chocoholic, I keep telling you to use y = mx + b. You can put it in standard form later.

OpenStudy (richyw):

just put it into slope intercept form and then rearrange it to standard form

OpenStudy (richyw):

you don't have to do this but it's much easier

OpenStudy (anonymous):

How do you put slope intercept form

OpenStudy (saifoo.khan):

See the last question we did @777chocoholic

OpenStudy (anonymous):

okaaaaaaaaay

OpenStudy (saifoo.khan):

And everyone else, don't help him.

hero (hero):

Sorry @saifoo.khan, I come here to help people.

hero (hero):

All I used was y = mx + b. No slope formulas or anything extra.

OpenStudy (saifoo.khan):

While people aren't keen to learn, which is wrong.

OpenStudy (anonymous):

qua qua

hero (hero):

Actually, that's not fully in standard form so you can finish it @777chocoholic

OpenStudy (anonymous):

3/5x + y = -4 , it doent let me input it its __=______

OpenStudy (saifoo.khan):

Dafuq does that supposed to mean? lol

hero (hero):

In general, the standard form of a line is \(Ax + By = C\)

OpenStudy (anonymous):

3x² + 5x - 20 = 0 ?

hero (hero):

Is that a new problem?

OpenStudy (anonymous):

no

hero (hero):

That's not what I posted for standard form.

hero (hero):

I'm sure you can clearly see what I posted. I even posted the general form

hero (hero):

Oh, I see. I made a mistake when solving for m.

hero (hero):

Too bad people are too lazy to back me up.

OpenStudy (anonymous):

haha

hero (hero):

-4 = 0m + b 1 = 5m + b b = -4 1 = 5m -4 1 + 4 = 5m 5 = 5m 1 = m Now that you have m and b, substitute it back into y = mx + b y = x - 4 Subtract y from both sides; add 4 to both sides to get: x - y = 4

hero (hero):

There we go. That should be it.

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

standard form?

hero (hero):

Yes, that is standard form for a line.

hero (hero):

It says standard form for the equation of a line

OpenStudy (anonymous):

4=x-y

OpenStudy (saifoo.khan):

x - y = 4

OpenStudy (anonymous):

?

hero (hero):

Standard form for the Equation of a line is \(Ax + By = C\) In this case, A = 1, B = -1, C = 4

OpenStudy (anonymous):

im in a fraking quiz and i got much time for each problem my gosh

OpenStudy (anonymous):

i jst need the answer pleaze

OpenStudy (saifoo.khan):

pleaZe

hero (hero):

Yes, that's exactly what I got @saifoo.khan. I made a mistake earlier, which is corrected.

OpenStudy (richyw):

that's plagiarism

OpenStudy (anonymous):

shut up

OpenStudy (saifoo.khan):

Okay, so im getting death threats in my inbox messages.

OpenStudy (anonymous):

by who???

hero (hero):

I basically told @saifoo.khan to stop posting stuff like ~pot, ~weed, and ~alcohol.

hero (hero):

That's what he calls "death threats"

OpenStudy (anonymous):

hero your a teachers pet

OpenStudy (saifoo.khan):

LMFAO no comments.

hero (hero):

I don't have a teacher and I don't own any pets.

hero (hero):

So, no, I'm not a teacher's pet, sorry.

OpenStudy (anonymous):

N3RD

OpenStudy (saifoo.khan):

looooool.

hero (hero):

I try to help @777chocoholic out, and this is the thanks I get.

OpenStudy (anonymous):

all i need is a feaking aswers im in a timed F***ing quiz

hero (hero):

Who's fault is that? Not mine. You shouldn't be on OS while taking a quiz anyway.

OpenStudy (saifoo.khan):

:'(

hero (hero):

This is not the place to get answers. Sorry.

OpenStudy (richyw):

This isn't a chat room. This also isn't the place to cheat on a quiz. Use the "chat" to chat so notifications quit popping up. Don't call us nerds and then expect us to help you cheat. The nerds didn't get to where we are by cheating.

OpenStudy (anonymous):

so 94y